# AdaSTaR: Adaptive Data Sampling for Training Self-Taught Reasoners

### v1

@@ 1 Introduction
Language models (LMs) are demonstrating remarkable emergent abilities across diverse cognitive tasks such as mathematical reasoning (Yao et al., 2023; Chen and Li, 2024; Brown et al., 2024), code generation (Sun et al., 2024; Research, 2025), and commonsense reasoning (Team, 2023; Google, 2023). Although LMs acquire foundational reasoning capabilities from large-scale pre-training and supervised finetuning (SFT), generating high-quality, explicit reasoning steps, often called Chains-of-Thought (CoT) (Wei et al., 2022a, 2023, b; Wang et al., 2023a), typically requires costly human annotation (Lightman et al., 2024; Havrilla et al., 2024; Zelikman et al., 2024). Creating such datasets is expensive and scales poorly, presenting a critical bottleneck as tasks increase in complexity. This challenge motivates the development of methods that improve LM reasoning without relying on extensive human annotation.
@@
We propose Adaptive STaR (AdaSTaR), a novel method that integrates adaptive sampling into the STaR training loop. AdaSTaR implements two core intuitions: (1) Adaptive Sampling for Diversity: prioritizing under-trained examples to ensure balanced learning; and (2) Adaptive Sampling for Curriculum: regularizing the system to sample easier data when the model is weaker early on. We empirically validate the effectiveness and efficiency of AdaSTaR through experiments across six reasoning datasets and an extensive list of baselines. AdaSTaR consistently improves both performance and computational efficiency. Remarkably, AdaSTaR not only achieves the highest test accuracy across all 6/6 benchmarks, but also simultaneously reduces the required training compute (PFLOPs) by an average of 58.6% compared to the strongest accuracy baseline (see Fig. 1). These performance and efficiency gains generalize to other pre-trained LMs and larger model size which we discuss further later.
@@
Although many works build on STaR, none, to our knowledge, target improving efficiency. Subsequent works improve performance at significant compute cost; AdaSTaR is complementary, improving scalability and accessibility. V-STaR (Hosseini et al., 2024) adds a verifier LM to improve inference-time performance through best-of-$N$ sampling (Snell et al., 2025). Iterative Reasoning Preference Optimization (Pang et al., 2024) incorporates a Direct Preference Optimization (Rafailov et al., 2023) term in its objective. To curate preference pairs, it increases CoT samples from $K=2$ in STaR to $K=30$. B-STaR (Zeng et al., 2025) enhances LM exploration for more diverse reasoning, and trains a separate process reward model (Uesato et al., 2022; Lightman et al., 2024) for finer-grained verification. Lean-STaR (Lin et al., 2025) employs the Lean theorem prover (De Moura et al., 2015) and a frontier LM (GPT-4) to extend STaR to mathematical theorem proving.
Reinforcement Learning (RL) offers a parallel approach to enhance LM reasoning, also leveraging an iterative process. RL’s reward-based objective often yields long-CoTs (Shao et al., 2024; DeepSeek-AI, 2025; Liu et al., 2025b; Yu et al., 2025b; Sui et al., 2025; Team, 2025; Liu et al., 2025a; Yu et al., 2025a), unlike the short-CoTs (NVIDIA, 2025) typical of STaR-style SFT. While the significantly larger token generation size of RL-based long-CoTs result in top performers, integrating STaR’s SFT remain a salient part of the training pipeline (Sui et al., 2025). For instance, Kimi k1.5 (Team, 2025), a representative reasoning model, utilizes STaR to expand their primary SFT dataset. To address the difficult, mixed-language, and overly long CoTs, DeepSeek-R1 (DeepSeek-AI, 2025) and Kimi k1.5 incorporate a STaR stage. Finally, DeepSeek-GRM (Liu et al., 2025b), a generalist reward model, also adopts a modified STaR as its training’s first stage. While these RL-based advancements are significant, our work concentrates on enhancing the efficiency and effectiveness of the STaR stage.
@@ 2.1 Preliminary: Self-Taught Reasoner (STaR) and its Variants
Let $\pi_{\theta}^{t}$ denote a LM (Vaswani et al., 2017) parameterized by $\theta$ at iteration $t$. We are given a supervised dataset $\mathcal{D}=\{\langle x_{i},y_{i}\rangle\}_{i=1}^{N}$. Following Wei et al. (2022b), each task is represented as $\langle x,c,y\rangle$, where $x\in\mathcal{X}$ is the query (input), $c\in\mathcal{C}$ is the CoT reasoning step(s), and $y\in\mathcal{Y}$ is the final answer. Since ground-truth CoTs $\mathcal{C}$ are unavailable, STaR aims to generate appropriate $c$ to improve generalization. To achieve this, $\pi_{\theta}^{t}$ generates $\langle\hat{c}_{i},\hat{y}_{i}\rangle$ conditioned on fixed few-shot CoT exemplars $e=\{\langle x_{e},c_{e},y_{e}\rangle\}_{e=1}^{E}$. However, as no ground truth $c_{i}$ is available, we require sampling and verification. Given the supervised dataset, a rule-based verifier defines a reward signal $r:=\mathbb{I}(y_{i}=\hat{y}_{i})$, where $\mathbb{I}(\cdot)$ is the indicator function.
$K\in\mathbb{N}$ number of CoT traces are sampled as follows (Fig. 2, blue). For the first $k\in[K]$, each observation $i$ is sampled once via $\langle\hat{c}_{i},\hat{y}_{i}\rangle\leftarrow\pi_{\theta}^{t}(e,x_{i})$. If $r=1$, it is accepted, and if $r=0$, it is resampled using rationalization (Zelikman et al., 2022): $\pi_{\theta}^{t}(e,x_{i}\oplus y_{i})$, where the ground truth $y_{i}$ is concatenated. In some extensions of STaR, $K>2$ samples are drawn without rationalization (Singh et al., 2024; Hosseini et al., 2024; Pang et al., 2024; Zeng et al., 2025; Lin et al., 2025).
@@ 2.2 Motivation: Need for Adaptive Data Sampling
[…]
@@ 12                   $w_{i}^{tmp}\leftarrow\frac{k-1}{k}w_{i}^{tmp}+\frac{1}{k}\mathbb{I}[\hat{y}_{i}=y_{i}]$;
18       $\alpha,\pi^{t+1}_{\theta}\leftarrow\texttt{Train}(\pi^{t}_{\theta},\mathcal{D}^{t}_{+})$ ;
@@ 21             $\tilde{t}_{i}\leftarrow t$, $w_{i}\leftarrow w^{tmp}_{i}$;
At each iteration $t$, we want our model to maximize $p_{i}^{t}:=\mathbb{P}(y_{i}=\hat{y}_{i}\leftarrow\pi_{\theta}^{t}(x_{i}))$ for all $i$’s. As the model is fitted with likelihood maximization (Fisher, 1922), we can expect $p_{i}^{t+1}\geq p_{i}^{t}$ for any $i$ sampled at iteration $t$. It is therefore sensible to prioritize observations with the lowest $p_{i}^{t}$ values, as they require more sampling and can be interpreted as more difficult at iteration $t$. It now remains to approximate $p_{i}^{t}$. A direct Monte Carlo estimate with $K$ samples gives $p_{i}^{t}\approx\hat{p}_{i}^{t}:=\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}[y_{i}=\hat{y}_{i}\leftarrow\pi_{\theta}^{t}(x_{i})]$. However, computing this for every $i$ at every iteration $t$ requires $K$ forward passes per observation, which is computationally prohibitive. Instead, we reuse the most recent estimate $\hat{p}_{i}^{{\tilde{t}_{i}}}$. The win static computation at ${\tilde{t}_{i}}$ induces no (run-time) compute overhead as the $K$ samples are an inherent part of the existing STaR system. Recalling that ${\tilde{t}_{i}}<t$ refers to the last iteration in which $i$ was sampled, $\hat{p}_{i}^{{\tilde{t}_{i}}}$ is the most recently available approximation to $\hat{p}_{i}^{t}$. Moreover, as we are priority-sorting w.r.t. ${\tilde{t}_{i}}$, we can expect that $t-{\tilde{t}_{i}}$ is not too large, i.e., $\hat{p}_{i}^{t}\approx\hat{p}_{i}^{{\tilde{t}_{i}}}$.
@@
As input, AdaSTaR takes the original dataset $\mathcal{D}$, base model $\pi^{t=1}_{\theta}$, and $n$-Shot CoT examplar $e$. For all observations, the statistics are initialized to 00 (lines 1, 2). In line 3, we utilize the Hierarchical Min Heap HieMinHeap (Cormen et al., 2022) to order the observations via the two statistics as follows: for two observations $i,j\in\texttt{HieMinHeap}(\cdot,\tilde{t},w)$,
$\underbrace{i\succ j}_{\text{$i$ is peeked/popped before $j$}}\Longleftrightarrow\underbrace{\tilde{t}_{i}<\tilde{t}_{j}}_{\text{$i$ is lastsampled before $j$}}\vee\ \underbrace{\left(\tilde{t}_{i}=\tilde{t}_{j}\wedge w_{i}<w_{j}\right)}_{\begin{subarray}{c}\text{$i$ and $j$ are last sampled at the}\\ \text{same $t$, but $i$ is more difficult}\end{subarray}}.$ (1)
@@ 3.2 Adaptive Data Sampling for Curriculum
We use the training accuracy $\alpha\in[0,1]$ from the current iteration $t$ as a proxy for model strength (Alg. 1, line 15). When $\alpha$ is low (indicating a weaker model), a relatively easier mix of observations should be prioritized for subsequent sampling. This regularization is automatically phased out as $\alpha$ increases with training. Similar to tracking $\tilde{t}_{i}$ and $w_{i}$, using $\alpha$ introduces no additional computational overhead, as the training step (which yields $\alpha$) is integral to the system. This explains our choice over, for instance, validation or test set accuracy; while potentially a more robust measures of generalization, these would require additional inference passes not intrinsic to the STaR loop.
@@ 4.1 Experimental Protocol
We conduct our main experiments with Llama 3.2 3B (Team, 2024b). We also evaluate using Qwen 2.5 3B (Team, 2024c) and Gemma 7B (Team, 2024a) to demonstrate the generality of our method across different model families. All base models are pre-trained-only models. For fairness, we optimize hyperparameters using the original STaR and apply them consistently across all methods. Further experimental details are provided in Appendix § E.
@@
We use two evaluation metrics: Test Set Accuracy (Acc.) and Floating Point Operations (FLOPs). The corresponding early-stopped (Caruana et al., 2000) epoch (e) and iteration (it) for vanilla SFT and STaR-based approaches, respectively are reported. All methods are given an equal and large compute budget to ensure that the peak value is obtained via early-stopping. For reproducibility, we evaluate accuracy using zero-shot greedy decoding unless stated otherwise. We use FLOPs as our computational cost metric as memory usage remains approximately constant across methods. FLOPs are computed empirically following Kaplan et al. (2020), Sardana et al. (2024).
@@ 4.2 Results
Table 1: Empirical results where Test Set Accuracy (%, $\uparrow$) is reported under zero-shot greedy decoding, excluding the 5-SC evaluation. Total training costs are reported in Peta FLOPs ($\downarrow$). Best Acc. and PFLOPs is bolded, and second best is underlined in each section (excluding SFT). In ( red) we quantify percent PFLOPs reduction against the highest accuracy baseline.
@@
AdaSTaR (ours) 73.8 10 it 174.4 ( $\downarrow$ 72.7%) 78.0 $20$ it 779.3 ( $\downarrow$ 32.7%) 95.6 23 it 3610.0 ( $\downarrow$ 19.2%)
@@
Comparing our AdaSTaR to baselines, AdaSTaR performs best in terms of accuracy in 6 of 6 benchmarks relative to 10 baselines, all while reducing training FLOPS by 58.6% on average (minimum of 19.2% to a maximum of 93.7%) relative to the strongest accuracy-achieving baseline. If there are numerous tied best baselines, we use the lowest PFLOPs to be conservative. Finally, for an intuitive visual understanding of our HieMinHeap, we provide empirical visualizations in Appendix § G.
@@ Results.
AdaSTaR wo. $w_{i}$ does indeed, on average, reduce SD, but fails to meaningfully improve test accuracy. Therefore, we can conclude that leveraging $w_{i}$ to induce sampling more challenging observations within tied $\tilde{t}_{i}$ is a salient part of AdaSTaR. We can decompose the rise in training diversity by quantifying the fall in SD throughout STaR-Acc $\rightarrow$ AdaSTaR wo. $w_{i}$rightarrow AdaSTaR: 1.72 $\rightarrow$ 1.65 $\rightarrow$ 1.45. AdaSTaR-PF fails to reduce SD, as it aggressively samples challenging observations ($\downarrow w_{i}$), resulting in frequent resampling of difficult examples. It also results in worsened test accuracy, likely due to poorer CoT quality (see § 3.2).

### v3

@@ 1 Introduction
Language models (LMs) are demonstrating remarkable emergent abilities across diverse cognitive tasks such as mathematical reasoning (Yao et al., 2023; Chen and Li, 2024; Brown et al., 2024), code generation (Sun et al., 2024; Research, 2025), and commonsense reasoning (Qwen Team, 2023; Google, 2023). Although LMs acquire foundational reasoning capabilities from large-scale pre-training and supervised finetuning (SFT), generating high-quality, explicit reasoning steps, often called Chains-of-Thought (CoT) (Wei et al., 2022a, 2023, b; Wang et al., 2023a), typically requires costly human annotation (Lightman et al., 2024; Havrilla et al., 2024; Zelikman et al., 2024). Creating such datasets is expensive and scales poorly, presenting a critical bottleneck as tasks increase in complexity. This challenge motivates the development of methods that improve LM reasoning without relying on extensive human annotation.
@@
We propose Adaptive STaR (AdaSTaR), a novel method that integrates adaptive sampling into the STaR training loop. AdaSTaR implements two core intuitions: (1) Adaptive Sampling for Diversity: prioritizing under-trained examples to ensure balanced learning; and (2) Adaptive Sampling for Curriculum: regularizing the system to sample easier data when the model is weaker early on. We empirically validate the effectiveness and efficiency of AdaSTaR through experiments across six reasoning datasets and an extensive list of baselines. AdaSTaR consistently improves both performance and computational efficiency. Remarkably, AdaSTaR not only achieves the highest test accuracy across all 6/6 benchmarks, but also simultaneously reduces the required training compute (FLOPs) by an average of 58.6% compared to the strongest accuracy baseline (see Fig. 1). These performance and efficiency gains generalize to other pre-trained LMs and larger model size which we discuss further later.
@@
Although many works build on STaR, none, to our knowledge, target improving efficiency. Subsequent works improve performance at significant compute cost; AdaSTaR is complementary, improving scalability and accessibility. V-STaR (Hosseini et al., 2024) adds a verifier LM to improve inference-time performance through best-of-$N$ sampling (Snell et al., 2025). Iterative Reasoning Preference Optimization (Pang et al., 2024) incorporates a Direct Preference Optimization (Rafailov et al., 2023) term in its objective: to curate preference pairs, it increases CoT samples from $K=2$ in STaR to $K=30$. B-STaR (Zeng et al., 2025) enhances LM exploration for more diverse reasoning, and trains a separate process reward model (Uesato et al., 2022; Lightman et al., 2024) for finer-grained verification. Lean-STaR (Lin et al., 2025) employs the Lean theorem prover (De Moura et al., 2015) and a frontier LM (GPT-4) to extend STaR to mathematical theorem proving.
Reinforcement Learning (RL) offers a parallel approach to enhance LM reasoning, also leveraging an iterative process. RL’s reward-based objective often yields long-CoTs (Shao et al., 2024; DeepSeek-AI, 2025; Liu et al., 2025b; Yu et al., 2025b; Sui et al., 2025; Kimi Team, 2025; Liu et al., 2025a; Yu et al., 2025a), unlike the short-CoTs (NVIDIA, 2025) typical of STaR-style SFT. While the significantly larger token generation size of RL-based long-CoTs result in top performers, integrating STaR’s SFT remain a salient part of the training pipeline (Sui et al., 2025). For instance, Kimi k1.5 (Kimi Team, 2025), a representative reasoning model, utilizes STaR to expand their primary SFT dataset. To address the difficult, mixed-language, and overly long CoTs, DeepSeek-R1 (DeepSeek-AI, 2025) and Kimi k1.5 incorporate a STaR stage. Finally, DeepSeek-GRM (Liu et al., 2025b), a generalist reward model, also adopts a modified STaR as its training’s first stage. While these RL-based advancements are significant, our work concentrates on enhancing the STaR stage.
@@ 2.1 Preliminary: Self-Taught Reasoner (STaR) and its Variants
Let $\pi_{\theta}^{t}$ denote a LM (Vaswani et al., 2017) parameterized by $\theta$ at iteration $t$. We are given a supervised dataset $\mathcal{D}=\{\langle x_{i},y_{i}\rangle\}_{i=1}^{N}$. Following Wei et al. (2022b), each task is represented as $\langle x,c,y\rangle$, where $x\in\mathcal{X}$ is the query (input), $c\in\mathcal{C}$ is the CoT reasoning step(s), and $y\in\mathcal{Y}$ is the final answer. Since ground-truth CoTs $\mathcal{C}$ are unavailable, STaR aims to generate appropriate $c$ to improve generalization. To achieve this, $\pi_{\theta}^{t}$ generates $\langle\hat{c}_{i},\hat{y}_{i}\rangle$ conditioned on fixed few-shot CoT exemplars $e=\{\langle x_{\epsilon},c_{\epsilon},y_{\epsilon}\rangle\}_{\epsilon=1}^{E}$. However, as no ground truth $c_{i}$ is available, we require sampling and verification. Given the supervised dataset, a rule-based verifier defines a reward signal $r:=\mathbb{I}(y_{i}=\hat{y}_{i})$, where $\mathbb{I}(\cdot)$ is the indicator function.
Let $K\in\mathbb{N}$ denote the number of CoT traces sampled as follows (Fig. 2, blue). For the first $k\in\{1,2,\cdots,K\}$, each observation $i$ is sampled once via $\langle\hat{c}_{i},\hat{y}_{i}\rangle\leftarrow\pi_{\theta}^{t}(e,x_{i})$. If $r=1$, it is accepted, and if $r=0$, it is resampled using rationalization (Zelikman et al., 2022): $\pi_{\theta}^{t}(e,x_{i}\oplus y_{i})$, where the ground truth $y_{i}$ is concatenated. In some extensions of STaR, $K>2$ samples are drawn without rationalization (Singh et al., 2024; Hosseini et al., 2024; Pang et al., 2024; Zeng et al., 2025; Lin et al., 2025).
@@ 2.2 Motivation: Need for Adaptive Data Sampling
Following Wei et al. (2025) and Lee et al. (2025), we leverage the strongest available teacher model (Ho et al., 2023) (GPT 4o) to annotate false positives. We compare a method encouraging diversity by sampling challenging observations (AdaD) against its baseline, STaR-ACC. AdaD and STaR-Acc are formally introduced in § 3 and 4.1. For each method, 100 observations are randomly sampled (without replacement) from $\mathcal{D}_{+}^{t}$ for CoT annotation. The precise iteration $t$ for both methods is chosen by taking $t:=\textnormal{min}(\textnormal{BestIter}(\texttt{AdaD}),\textnormal{BestIter}(\texttt{STaR-Acc}))$, where $\textnormal{BestIter}(\cdot)$ is the early-stopped iteration. Further details and a qualitative example are provided in Appendix § C.
@@ 12       $w_{i}^{tmp}\leftarrow\frac{k-1}{k}w_{i}^{tmp}+\frac{1}{k}\mathbb{I}[\hat{y}_{i}=y_{i}]$;
18 ${\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\alpha},\pi^{t+1}_{\theta}\leftarrow\texttt{Train}(\pi^{t}_{\theta},\mathcal{D}^{t}_{+})$ ;
@@ 21    $\tilde{t}_{i}\leftarrow t$, $w_{i}\leftarrow w^{tmp}_{i}$;
At each iteration $t$, we want our model to maximize $p_{i}^{t}:=\mathbb{P}(y_{i}=\hat{y}_{i}\leftarrow\pi_{\theta}^{t}(x_{i}))$ for all $i$’s. As the model is fitted with likelihood maximization (Fisher, 1922), we can expect $p_{i}^{t+1}\geq p_{i}^{t}$ for any $i$ sampled at iteration $t$. It is therefore sensible to prioritize observations with the lowest $p_{i}^{t}$ values, as they require more sampling and can be interpreted as more difficult at iteration $t$. It now remains to approximate $p_{i}^{t}$. A direct Monte Carlo estimate with $K$ samples gives $p_{i}^{t}\approx\hat{p}_{i}^{t}:=\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}[y_{i}=\hat{y}_{i}\leftarrow\pi_{\theta}^{t}(x_{i})]$. However, computing this for every $i$ at every iteration $t$ requires $K$ forward passes per observation, which is computationally prohibitive. Instead, we reuse the most recent estimate $\hat{p}_{i}^{{\tilde{t}_{i}}}$. The win static computation at $\tilde{t}_{i}$} induces no (run-time) compute overhead as the $K$ samples are an inherent part of the existing STaR system. Recalling that ${\tilde{t}_{i}}<t$ refers to the last iteration in which $i$ was sampled, $\hat{p}_{i}^{{\tilde{t}_{i}}}$ is the most recently available approximation to $\hat{p}_{i}^{t}$. Moreover, as we are priority-sorting with respect to $\tilde{t}_{i}$}, we can expect that $t-{\tilde{t}_{i}}$ is not too large, i.e., $\hat{p}_{i}^{t}\approx\hat{p}_{i}^{{\tilde{t}_{i}}}$.
@@
As input, AdaSTaR takes the original dataset $\mathcal{D}$, base model $\pi^{t=1}_{\theta}$, and $n$-Shot CoT examplar $e$. For all observations, the statistics are initialized to 0 (lines 1, 2). In line 3, we utilize Cormen et al. (2022)’s Hierarchical Min Heap HieMinHeap to order the observations via the two statistics as follows: for two observations $i,j\in\texttt{HieMinHeap}(\cdot,\tilde{t},w)$,
[…]
@@ 3.2 Adaptive Data Sampling for Curriculum
We use the training accuracy $\alpha\in[0,1]$ from the current iteration $t$ as a proxy for model strength (Alg. 1, line 15). When $\alpha$ is low (indicating a weaker model), a relatively easier mix of observations should be prioritized for subsequent sampling. This regularization is automatically phased out as $\alpha$ increases with training. Similar to tracking $\tilde{t}_{i}$ and $w_{i}$, using $\alpha$ introduces no additional computational overhead, as the training step (which yields $\alpha$) is integral to the system. This explains our choice over, for instance validation set accuracy (not used in final evaluation); while potentially a more robust measures of generalization, these would require additional inference passes not intrinsic to the STaR loop.
@@ 4.1 Experimental Protocol
We conduct our main experiments with Llama 3.2 3B (Llama Team, 2024). We also evaluate using Qwen 2.5 3B (Qwen Team, 2024) and Gemma 7B (Gemma Team, 2024) to demonstrate the generality of our method across different model families. All base models are pre-trained-only models. For fairness, we optimize hyperparameters using the original STaR and apply them consistently across all methods. Further experimental details are provided in Appendix § E.
@@
We use two evaluation metrics: Test Set Accuracy (Acc.) and Floating Point Operations (FLOPs). The corresponding early-stopped (Caruana et al., 2000) epoch (e) and iteration (it) for vanilla SFT and STaR-based approaches, respectively are reported. All methods are given an equal and large compute budget to ensure that the peak value is obtained via early-stopping. For reproducibility, we evaluate accuracy using zero-shot greedy decoding unless stated otherwise. We use FLOPs as our computational cost metric as memory usage remains approximately constant across methods. FLOPs are computed empirically following the method used by Kaplan et al. (2020), Sardana et al. (2024).
@@ 4.2 Results
Table 1: Empirical results where Test Set Accuracy (%, $\uparrow$) is reported under zero-shot greedy decoding, excluding the 5-SC evaluation. Total training costs are reported in Peta FLOPs (PFLOPs, $\downarrow$). Best Acc. and PFLOPs is bolded, and second best is underlined in each section (excluding SFT). In ( red) we quantify percent PFLOPs reduction against the highest accuracy baseline.
@@
AdaSTaR (ours) 73.8 10 it 174.4 ( $\downarrow$ 72.7%) 78.0 2020 it 779.3 ( $\downarrow$ 32.7%) 95.6 23 it 3610.0 ( $\downarrow$ 19.2%)
@@
Comparing our AdaSTaR to baselines, AdaSTaR performs best in terms of accuracy in 6 of 6 benchmarks relative to 10 baselines, all while reducing training FLOPS by a mean of 58.6% (minimum of 19.2% to a maximum of 93.7%) relative to the strongest accuracy-achieving baseline. If there are numerous tied best baselines, we use the lowest PFLOPs to be conservative. Finally, for an intuitive visual understanding of our HieMinHeap, we provide empirical visualizations in Appendix § G.
@@ Results.
AdaSTaR wo. $w_{i}$ does indeed, on average, reduce SD, but fails to meaningfully improve test accuracy. Therefore, we can conclude that leveraging $w_{i}$ to induce sampling more challenging observations within tied $\tilde{t}_{i}$ is a salient part of AdaSTaR. We can decompose the rise in training diversity by quantifying the fall in SD throughout STaR-Acc $\rightarrow$ AdaSTaR wo. $w_{i}$ $\rightarrow$ AdaSTaR: 1.72 $\rightarrow$ 1.65 $\rightarrow$ 1.45. AdaSTaR-PF fails to reduce SD, as it aggressively samples challenging observations ($\downarrow w_{i}$), resulting in frequent resampling of difficult examples. It also results in worsened test accuracy, likely due to poorer CoT quality (see § 3.2).